help with linear algebra hw problem I get incorrect
Diagonalize the following matrix, if possible. А 3 3 5 1
Answer #1

3 3 (AxI=0 5 3-1 3 |=0 (3-X) (1-1) - 150 3 + x 2 - 4 - 15=0 → 1-41 -12 = 0 - 68+2x - 12=0 = (1-6) +2 (1-6)=0 → (1-6) (4+2) =(A-GI) x = 0 - 15:00 a-ng = x=xq So = ng (3) = {[:)] E6 és a basis for the eigen space -3 [ 2 0 6 Now AP - 3 3 [ 5 5 6 - 6 6